### v1

# SegSTRONG-C: Segmenting Surgical Tools Robustly On Non-adversarial Generated Corruptions – An EndoVis’24 Challenge

Abstract
Accurate segmentation of tools in robot-assisted surgery is critical for machine perception, as it facilitates numerous downstream tasks including augmented reality feedback. While current feed-forward neural network-based methods exhibit excellent segmentation performance under ideal conditions, these models have proven susceptible to even minor corruptions, significantly impairing the model’s performance. This vulnerability is especially problematic in surgical settings where predictions might be used to inform high-stakes decisions.
To better understand model behavior under non-adversarial corruptions, prior work has explored introducing artificial corruptions, like Gaussian noise or contrast perturbation to test set images, to assess model robustness. However, these corruptions are either not photo-realistic or model/task agnostic. Thus, these investigations provide limited insights into model deterioration under realistic surgical corruptions.
To address this limitation, we introduce the SegSTRONG-C challenge that aims to promote the development of algorithms robust to unforeseen but plausible image corruptions of surgery, like smoke, bleeding, and low brightness. We collect and release corruption-free mock endoscopic video sequences for the challenge participants to train their algorithms and benchmark them on video sequences with photo-realistic non-adversarial corruptions for a binary robot tool segmentation task. The training set consists of $11$ corruption-free mock endoscopic surgery sequences, in total of $6600$ frames, and corresponding segmentation masks representing surgical tools. The validation set consists of $3$ video sequences, each with and without corrupted background, in total $3600$ frames. The test set consists of $3$ sequences with non-adversarial corruptions (smoke, bleeding, and low brightness), in total $5400$ frames. This new benchmark will allow us to carefully study neural network robustness to non-adversarial corruptions of surgery, thus constituting an important first step towards more robust models for surgical computer vision. In this paper, we describe the data collection and annotation protocol, baseline evaluations of established segmentation models, and data augmentation-based techniques to enhance model robustness.
Background & Summary
Background:
Robot tool segmentation is a fundamental machine perception task in robot-assisted surgery, enabling numerous downstream tasks. Feed-forward networks (FFN) [1, 2, 3, 4, 5, 6, 7, 8] have dominated this task since deep learning’s emergence and the introduction of public surgical scenes datasets through EndoVis [9, 10] challenges. Methods that achieve state-of-the-art performance on general vision benchmarks [11, 12] also exhibit excellent performance on benchmarks from the surgical domain as the test cases are usually sampled from the same distribution as the training cases and contain minimal corruptions. However, FFNs’ susceptibility to even minor corruptions can significantly impair their performance. This vulnerability is especially critical in high-stakes surgical decisions and poses a significant barrier to clinical translation [13].
Current datasets and data collection/synthesis strategies used to evaluate non-adversarial robustness represented by ImageNet-C [14] are mainly generated synthetically [15]. Thus, their ability to represent corruption likely to occur in surgical scenes remains unknown. To generate test cases that can assess algorithms’ robustness against common complications observed during the surgery, such as smoke, bleeding, and low brightness, Ding et al. [16, 17] utilized the accurate replay functionality of the da Vinci Research Kit(dVRK) [18] to add real-world corruptions to the same trajectory. In this work, we adopt the same data collection pipeline from  [16] and refine it by incorporating the segment anything model (SAM) for more precise and efficient annotations to collect the dataset to assess the algorithms’ robustness against common complications in the surgical scenario.
There are methods proposed to address robustness under non-adversarial corruption with different tactics [15]. Architecture-based tactics address the robustness by developing full architectures [19, 20, 21] or motifs [22, 23, 24, 25]. Data augmentation [26, 27, 28, 29, 30] tactics aim to encourage models to be invariant to non-causal features by increasing the diversity of the training dataset. Optimization-based robustness tactics [31, 32, 33, 34, 35] modify learning objectives to improve resilience to corruption. Auxiliary data(e.g. kinematics in robotics surgery), when available, is used to improve the robustness [17, 16]. However, despite the efforts, a measurable gap remains [15].
We propose the SegSTRONG-C challenge to encourage more attention to the models’ robustness in the healthcare domain and test the robustness under non-adversarial corruptions that commonly happen in the surgical scene to provide more trustworthy assessment and analysis for future directions.
Summary:
With the collected dataset we propose the SegSTRONG-C challenge where participants are challenged to train their algorithms solely on uncorrupted sequences and achieve high performance on corrupted ones for the binary robot tool segmentation task. In this early version of the paper, we present the dataset for the SegSTRONG-C challenge along with its collection and annotation pipeline, the evaluation metrics for the challenge, and a baseline analysis with some segmentation networks and data augmentation methods as baseline models.
Dataset
The dataset consists of mock endoscopic video sequences with corresponding binary segmentation masks for the robot tool segmentation task. We mock the endoscopic scene with two patient-side manipulators (PSMs) and animal tissue backgrounds to ensure a photo-realistic appearance. We manually teleoperate the robot to generate the trajectory in free space. The binary segmentation masks serve as ground truth annotations for surgical tools. The dataset consists of $17$ sequences, collected from different robot and camera configurations and different robot trajectories. Each sequence is collected at $10$ frames per second and consists of $300$ uncorrupted frames for the left camera and $300$ uncorrupted frames for the right camera. We collect all corrupted versions (background, smoke, bleeding, and low brightness) for each sequence under the same configuration and trajectory. We provide $11$ sequences of the dataset with only uncorrupted frames as the train set. We provide $3$ video sequences with uncorrupted frames and corresponding frames with background corruption as the validation set. We use the final 3 sequences with certain corruptions(smoke, bleeding, and low brightness) as the test set. The models submitted by the challenge participants will be tested on this test set consisting of sequences with photo-realistic non-adversarial corruptions. The models will be tested on each corruption separately. As the challenge is currently active, we withhold the release of the test set. Table 1 shows the summary of the dataset. All corrupted versions for the training and validation will be released after the challenge submission. The example images are shown in Figure 1.
Table 1: Dataset summary. "300 + 300" means 300 frames for the left camera and 300 frames for the right camera. "-" means not provided during challenge.
Split Camera/robot Trajectory Non- Background Smoke Bleeding Low
configuration id id corruption brightness
Train set 3 1 300 + 300 - - - -
3 2 300 + 300 - - - -
4 3 300 + 300 - - - -
4 4 300 + 300 - - - -
4 5 300 + 300 - - - -
5 6 300 + 300 - - - -
5 7 300 + 300 - - - -
7 8 300 + 300 - - - -
7 9 300 + 300 - - - -
8 10 300 + 300 - - - -
8 11 300 + 300 - - - -
Validation set 1 12 300 + 300 300 + 300 - - -
1 13 300 + 300 300 + 300 - - -
1 14 300 + 300 300 + 300 - - -
Test set 9 15 - - 300 + 300 300 + 300 300 + 300
9 16 - - 300 + 300 300 + 300 300 + 300
9 17 - - 300 + 300 300 + 300 300 + 300
Figure 1: Example images for regular image and non-adversarial corruptions(background, smoke, bleeding, and low brightness.).
Data Collection
The data was acquired in the Robotorium of the Laboratory for Computational Sensing and Robotics (LCSR), Johns Hopkins University by surgical robotics experts familiar with operating the da Vinci robot via dVRK [18]. We use dVRK [18] as our robot platform with the endoscopic camera manufactured by SCHÖLLY and the image process unit manufactured by Ikegami as the perception part. We first adjust the camera/robot configuration and perform calibrations. For each configuration, we collect different trajectories by human operation. For each trajectory, we replay the recorded kinematics to reproduce the trajectory and record video sequences under different scenarios including non-corrupted and corrupted versions with background, smoke, bleeding, and low brightness. To generate non-adversarial corruptions for the same testing samples, we record the robot trajectory, replay the kinematics via dVRK, and manually add corruption for each replay. The background corruption is generated by changing the type of background tissue. The smoke corruption is mimicked by adding artificial fog via a fog machine. The bleeding corruption is mimicked by fake blood. The low brightness corruption is generated by turning down the camera light. The data collection pipeline is shown in the following steps:
  • Step1: Camera calibration and hand-eye calibration - We perform a camera calibration to get the intrinsics for the left and right cameras of the stereo and a hand-eye calibration to get the transformation from both cameras to the base frame of both PSM1 and PSM2.
  • Step2: Trajectory generation - We use the teleoperation feature of the dVRK to manipulate the PSMs to generate trajectories in free space and record the kinematics of the trajectory.
  • Step3: Trajectory replay and recording - We replay the same trajectories and record the videos at 10 fps with the same robot configuration under different scenarios, (1) pure dark background samples for ground truth generation, (2) uncorrupted samples. (3) non-adversarial corrupted samples.
Data annotation
The annotations are generated via a semi-automatic pipeline based on exclusively collected sequences with purely black backgrounds where the only salient areas in the images are the PSMs. We first automatically generate a segmentation mask for the PSMs via the segment anything model (SAM) [36] where the prompts are generated via traditional background extraction algorithm. We select experts who are familiar with the robot and segmentation task to be the annotators to verify and correct the automatically generated annotation. The annotation pipeline can be expressed in the following steps:
  • Step 1: Prompt generation - We use a traditional background extraction algorithm to generate rough masks for the foreground PSMs. The rough masks are converted to the prompt points and bounding boxes for SAM.
  • Step 2: Automatic generation - We input the prompt with the image to generate a fine mask for the robot tool. The first two steps are fully automatic.
  • Step 3: Failure case selection - Human annotators are involved in examining the generated mask from SAM and selecting the failure cases. We had two annotators examining the same sequences and using the union of their selections as the set of failure cases.
  • Step 4: Manual correction - The human annotators manually refine the selected failure cases from step 3 via SAM with human prompt input until satisfactory results are achieved. If the annotator had three attempts but did not get satisfactory results, they would draw the contour to annotate this sample manually.
Evaluation
We introduce how we quantitatively evaluate the performance of an algorithm and how we decide the ranking of multiple algorithms that participate in the challenge.
Metrics:
We use dice similarity coefficient (DSC) and normalized surface distance (NSD) averaged from different tolerances for the robot tool for multiple corruptions - low brightness, smoke, and blood. The DSC is a widely used metric in the field of medical image analysis. For a 2D binary class image segmentation task, DSC is defined as
$DSC=\frac{2TP}{2TP+FP+FN}$
, where $TP$ is the number of true positive pixels for the class, $FP$ is the number of false positives and $FN$ is the number of false negatives. NSD measures the overlap of two boundaries between the predicted and ground truth segmentation masks. A boundary pixel is counted as overlapping when the closest distance to the other boundary is less or equal to the specified tolerance. NSD for a 2D binary class image segmentation task is defined as:
$NSD=\frac{\left|S_{A}\cap\mathcal{B}_{B}^{(\tau)}\right|+\left|S_{B}\cap\mathcal{B}_{A}^{(\tau)}\right|}{\left|S_{A}\right|+\left|S_{B}\right|}$
, where $A$ and $B$ are the masks, $S_{A}$ and $S_{B}$ are boundaries, and $\mathcal{B}_{A}^{(\mathrm{\tau})}$ and $\mathcal{B}_{B}^{(\mathrm{\tau})}$ are the boundary of $A$ and $B$ with an extended border of width tolerance $\tau$. We calculate the average DSC and NSD over images. DSC is the standard metric for segmentation while NSD is complementary to DSC. DSC reveals performance more on the chunk area while NDS focuses on the boundary. We evaluate the performance only on the unseen domain to encourage participants to focus on the algorithm’s robustness.
Ranking:
We rank algorithms with total scores summed over 3 test domains - bleeding, smoke, and low brightness. Each Domain will have two ranks based on the DSC and NSD metrics. For ranking each domain, we will first sort the method based on the metrics over images. Then, we do a pairwise significance test for adjacent methods to test whether they have statistically significant differences in performance. If yes, we rank them according to the order, otherwise they have the same rank. The mean value indicates the overall performance, and the pairwise significance test can justify whether one method has a significant advantage over the other one. We give points in a decreasing arithmetic sequence according to the rank for each test domain. For example, Rank from 1 to 5 and get points from 5 to 1 if there are 5 participants in total. We only test on the unseen corrupted domains as we care about the overall performance of the model on the corrupted domain. Since we have more than one test domain, the summation of points given by rank is a reasonable way to have fair importance for each domain. Using rank and ignoring absolute performance difference is to avoid the result from one specific domain and metric dominating others. Since the goal for the participating team is to develop methods that are robust against all corruptions instead of algorithms achieving dominating performance on one domain but failing to generalize to others, The proposed overall ranking method better suits the goal of the challenge.
Baseline Analysis
In this section, we first briefly introduce the selected segmentation networks and data augmentation methods. Then, we present the testing results of these methods and perform a brief analysis of the baseline results.
Networks
U-Net
: U-Net[2] is originally developed for biomedical image segmentation. The architecture of U-Net consists of two paths: a contracting path and an expansive path. The contracting path is characterized by repeated convolutions and resembles a conventional convolutional network, followed by a rectified linear unit(RELU) and a max pooling operation. On the other hand, the expansive path enhances feature and spatial information through a series of up-convolutions and concatenation of features from the contracting path. Our implementation uses a 4-level vanilla U-Net as the baseline model.
DeepLabV3+
: The most recent version of the DeepLab family is DeepLabV3+ [3]. The architecture is changed to an encoder-decoder structure. The encoder part makes use of atrous separable convolution which separates the convolution process into a Depthwise Convolution and a Pointwise Convolution. Depthwise Convolution applies different kernels to each channel, and Pointwise Convolution combines the outputs across channels. The decoder part combines the low-level features and the upsampled output from the encoder to recover the spatial information by upscaling.
SegFormer
: SegFormer [7] is a semantic segmentation framework that combines a hierarchical transformer encoder with a multi-layer perceptron (MLP) decoder. The hierarchically structured transformer encoder avoids the interpolation of positional codes and outputs multi-scaled features. The lightweight MLP decoder combines both local and global attention to aggregate the multi-layer information.
SETR
: SETR [6] provides an alternative perspective to the encoder-decoder-based FCN architecture. It treats semantic segmentation as a sequence-to-sequence prediction task in which the author splits the image into patches, linearly projects each patch, adds positional embeddings, and feeds the sequence into the transformer encoder. Three different decoders are designed to perform pixel-level segmentation, and we implement all three: (1) Naive upsampling which projects the transformer feature to the dimension of class number and directly upsample to the desired image resolution, (2) Progressive upsampling which alternates convolution and upsampling operations instead of one-step upscaling, and (3) Multi-Level feature aggregation which takes feature representations from $M$ transformer layers uniformly, reshapes and aggregates the feature map top-down, fuses features from all channels and upsamples to the full resolution.
Augmentation
Figure 2: Example image augmentations
AutoAugment
: AutoAugment [30] searches for the best sequence of transformations that improves the performance of a model on a given dataset. We modify the Pytorch implementation of the AutoAugmentPolicy class and choose the policy learned on ImageNet.
Elastic
: Elastic transformation generates displacement vectors for each image pixel based on random offsets. It will transform the image given $\alpha$ and $\beta$, where $\alpha$ controls the strength and $\beta$ controls the smoothness of displacements. We modify the Pytorch implementation of the ElasticTransform class. Let $\alpha_{x}$ and $\alpha_{y}$ be the magnitude of displacements, $\sigma_{x}$ and $\sigma_{y}$ be the smoothness of displacements along $x$ and $y$ axis, and $h$ and $w$ be the height and width of the image. Kernel size along $x$ axis is: $k_{x}=\text{int}(8\sigma_{x}+1)$ thus half of the kernel size is $k_{x\text{\_half}}=\frac{1}{2}(k_{x}-1)$, and the displacement vector $d_{x}$ is first initialized as an array of random numbers between $-1$ and $1$. Each pixel is defined as the Gaussian blur of $d_{x}$ times $\alpha_{x}$ and divided by image height $h$. The 2D Gaussian kernel $h_{x}$ is defined as $h_{x}[m,n]=h_{1}[m]h_{1}[n]^{T}$ where $h_{1}$ is the normalized discrete 1D Gaussian kernel of length $k_{x}$. $d_{x}$, $d_{y}$ is defined as $d_{x}[m,n]=\frac{\alpha_{x}}{h}d_{x}[m,n]*h_{x}[m,n]$, $d_{y}[m,n]=\frac{\alpha_{y}}{w}d_{y}[m,n]*h_{y}[m,n]$. The complete displacement vector field is obtained by concatenating $d_{x}$ and $d_{y}$. The detailed derivations of $h_{1}$, $d_{x}$, and $d_{y}$ are in the equation:
$h_{1}[m]=\frac{\exp[(-\frac{1}{2}(\frac{-k_{x\text{\_half}}+m}{\sigma_{x}})^{2})]}{\sum_{n=-k_{x\text{\_half}}}^{k_{x\text{\_half}}}\exp[(-\frac{1}{2}(\frac{-k_{x\text{\_half}}+n}{\sigma_{x}})^{2})]}$
$d_{x}[m,n]=\frac{\alpha_{x}}{h}\sum_{j=-\infty}^{\infty}\sum_{i=-\infty}^{\infty}d_{x}[i,j]\cdot h_{x}[m-i,n-j]$
$d_{y}[m,n]=\frac{\alpha_{y}}{w}\sum_{j=-\infty}^{\infty}\sum_{i=-\infty}^{\infty}d_{y}[i,j]\cdot h_{y}[m-i,n-j]$
Projective
: Projective transformation transforms the image such that it is viewed from a different viewing perspective with a given probability. We modify the Pytorch implementation of the RandomPerspective class. The four new coordinates of the corners of the image are first randomly generated according to the input distortion scale. In perspective transform, each pixel $(x,y)$ in the original image is mapped to its new coordinate by the formula: $(u_{i},v_{i})=(\frac{ax_{i}+by_{i}+c}{gx_{i}+hy_{i}+1},\frac{dx_{i}+ey_{i}+f}{gx_{i}+hy_{i}+1})$ for $i=0,1,2,3$. In matrix form, it can be written as:
$\begin{bmatrix}x_{0}&y_{0}&1&0&0&0&-x_{0}u_{0}-y_{0}u_{0}\\ x_{1}&y_{1}&1&0&0&0&-x_{1}u_{1}-y_{1}u_{1}\\ x_{2}&y_{2}&1&0&0&0&-x_{2}u_{2}-y_{2}u_{2}\\ x_{3}&y_{3}&1&0&0&0&-x_{3}u_{3}-y_{3}u_{3}\\ 0&0&0&x_{0}&y_{0}&1&-x_{0}v_{0}-y_{0}v_{0}\\ 0&0&0&x_{1}&y_{1}&1&-x_{1}v_{1}-y_{1}v_{1}\\ 0&0&0&x_{2}&y_{2}&1&-x_{2}v_{2}-y_{2}v_{2}\\ 0&0&0&x_{3}&y_{3}&1&-x_{3}v_{3}-y_{3}v_{3}\end{bmatrix}\cdot\begin{bmatrix}a\\ b\\ c\\ d\\ e\\ f\\ g\\ h\end{bmatrix}=\begin{bmatrix}u_{0}\\ u_{1}\\ u_{2}\\ u_{3}\\ v_{0}\\ v_{1}\\ v_{2}\\ v_{3}\end{bmatrix}$
Solve this system to obtain the least-square solution for the coefficients, and map each pixel to its new position by the perspective transformation formula.
Results
We evaluate the performance of baseline models with and without data augmentation across different domains. While all models show reasonable accuracy under regular conditions, their effectiveness substantially deteriorates when tested in shifted domains such as in the presence of bleeding, low brightness, and smoke. The data augmentations applied, including AutoAugment, elastic, and projective transformations, generally failed to improve model performance in these domains. Notably, UNet combined with AutoAugment showed a marginal improvement under smoke conditions but this improvement is inconsistent across all domains. As shown in Figure 3, under regular conditions, the models adequately segmented tools from the background, but struggled in areas with appearance change and subtle contrast changes. Failures are predominantly under bleeding conditions due to the appearance (color) change of the shaft, leading to incorrect segmentation of the shaft. Under low brightness conditions, the darker the region is, the more likely a failure in segmentation will occur. The pervasive false positives under smoke conditions indicate that the models interpret the bright regions (smoke) as the foreground. These results suggest that all baseline models are highly dependent on adequate targets’ saliency to successfully distinguish between the tools and the background. They are also sensitive to any appearance change for the tool (bleeding) and background (smoke).
Table 2: Experiment results of baseline models across different domains.
Models NSD DSC
Bleeding Smoke Low Brightness Bleeding Smoke Low Brightness
DeepLabv3 [3] 0.5629 0.4637 0.4000 0.6896 0.6538 0.5352
Segformer [7] 0.5133 0.5266 0.4194 0.6802 0.6906 0.6145
SETR_PUP [6] 0.2531 0.3354 0.2599 0.4717 0.5848 0.4053
SETR_MLA [6] 0.2798 0.3374 0.1571 0.5358 0.6206 0.3385
SETR_Naive [6] 0.3312 0.4409 0.0092 0.5064 0.6657 0.0275
UNet [2] 0.5677 0.5084 0.4390 0.7052 0.6603 0.5750
UNet [2] + AutoAugment [30] 0.6654 0.8152 0.5344 0.7910 0.8895 0.6965
UNet [2] + Elastic 0.5622 0.5207 0.3931 0.6910 0.6583 0.5190
UNet [2] + Projective 0.5661 0.5702 0.4265 0.6903 0.6978 0.5792
SETR [6] + AutoAugment [30] 0.1172 0.0910 0.0773 0.1918 0.2157 0.0934
SETR [6] + Elastic 0.2836 0.3363 0.1654 0.5350 0.6192 0.3472
SETR [6]+ Projective 0.2240 0.3055 0.0552 0.4168 0.5766 0.1447
Conclusion
We present the SegSTRONG-C challenge, aiming to assess algorithms’ robustness to unforeseen yet plausible complications in surgical scenarios. To generate non-adversarial corruptions we use and refine the data collection and annotation pipeline from CaRTS [16] and collect 17 sequences for training and evaluation. We use DSC and NSD as quantitative metrics to evaluate algorithms’ performance on the testing set and propose a balanced ranking method to alleviate the influence of extreme performance on a single domain while emphasizing the statistical significance during ranking. The baseline analysis provides baseline performance for participants to refer to while providing a basic insight that the models’ success relies on adequate saliency and that all models are sensitive to appearance change.
Acknowledgement:
This Challenge is supported by the collaborative research agreement with the MultiScale Medical Robotics Center at The Chinese University of Hong Kong, Intuitive Surgical Inc, and Johns Hopkins University internal funds.
Figure 3: Baseline results. The white area means true positive, the red area means false negative, and the orange area means false positive.

### v2

@@
 Abstract
-Accurate segmentation of tools in robot-assisted surgery is critical for machine perception, as it facilitates numerous downstream tasks including augmented reality feedback.
-While current feed-forward neural network-based methods exhibit excellent segmentation performance under ideal conditions, these models have proven susceptible to even minor corruptions, significantly impairing the model’s performance.
-This vulnerability is especially problematic in surgical settings where predictions might be used to inform high-stakes decisions.
-To better understand model behavior under non-adversarial corruptions, prior work has explored introducing artificial corruptions, like Gaussian noise or contrast perturbation to test set images, to assess model robustness.
-However, these corruptions are either not photo-realistic or model/task agnostic.
-Thus, these investigations provide limited insights into model deterioration under realistic surgical corruptions.
-To address this limitation, we introduce the SegSTRONG-C challenge that aims to promote the development of algorithms robust to unforeseen but plausible image corruptions of surgery, like smoke, bleeding, and low brightness.
-We collect and release corruption-free mock endoscopic video sequences for the challenge participants to train their algorithms and benchmark them on video sequences with photo-realistic non-adversarial corruptions for a binary robot tool segmentation task.
-The training set consists of $11$ corruption-free mock endoscopic surgery sequences, in total of $6600$ frames, and corresponding segmentation masks representing surgical tools.
-The validation set consists of $3$ video sequences, each with and without corrupted background, in total $3600$ frames.
-The test set consists of $3$ sequences with non-adversarial corruptions (smoke, bleeding, and low brightness), in total $5400$ frames.
-This new benchmark will allow us to carefully study neural network robustness to non-adversarial corruptions of surgery, thus constituting an important first step towards more robust models for surgical computer vision.
-In this paper, we describe the data collection and annotation protocol, baseline evaluations of established segmentation models, and data augmentation-based techniques to enhance model robustness.
-Background & Summary
-Background:
-Robot tool segmentation is a fundamental machine perception task in robot-assisted surgery, enabling numerous downstream tasks.
-Feed-forward networks (FFN) [1, 2, 3, 4, 5, 6, 7, 8] have dominated this task since deep learning’s emergence and the introduction of public surgical scenes datasets through EndoVis [9, 10] challenges.
-Methods that achieve state-of-the-art performance on general vision benchmarks [11, 12] also exhibit excellent performance on benchmarks from the surgical domain as the test cases are usually sampled from the same distribution as the training cases and contain minimal corruptions.
-However, FFNs’ susceptibility to even minor corruptions can significantly impair their performance.
-This vulnerability is especially critical in high-stakes surgical decisions and poses a significant barrier to clinical translation [13].
-Current datasets and data collection/synthesis strategies used to evaluate non-adversarial robustness represented by ImageNet-C [14] are mainly generated synthetically [15].
-Thus, their ability to represent corruption likely to occur in surgical scenes remains unknown.
-To generate test cases that can assess algorithms’ robustness against common complications observed during the surgery, such as smoke, bleeding, and low brightness, Ding et al.
-[16, 17] utilized the accurate replay functionality of the da Vinci Research Kit(dVRK) [18] to add real-world corruptions to the same trajectory.
-In this work, we adopt the same data collection pipeline from  [16] and refine it by incorporating the segment anything model (SAM) for more precise and efficient annotations to collect the dataset to assess the algorithms’ robustness against common complications in the surgical scenario.
-There are methods proposed to address robustness under non-adversarial corruption with different tactics [15].
-Architecture-based tactics address the robustness by developing full architectures [19, 20, 21] or motifs [22, 23, 24, 25].
-Data augmentation [26, 27, 28, 29, 30] tactics aim to encourage models to be invariant to non-causal features by increasing the diversity of the training dataset.
-Optimization-based robustness tactics [31, 32, 33, 34, 35] modify learning objectives to improve resilience to corruption.
-Auxiliary data(e.g.
-kinematics in robotics surgery), when available, is used to improve the robustness [17, 16].
-However, despite the efforts, a measurable gap remains [15].
-We propose the SegSTRONG-C challenge to encourage more attention to the models’ robustness in the healthcare domain and test the robustness under non-adversarial corruptions that commonly happen in the surgical scene to provide more trustworthy assessment and analysis for future directions.
-Summary:
-With the collected dataset we propose the SegSTRONG-C challenge where participants are challenged to train their algorithms solely on uncorrupted sequences and achieve high performance on corrupted ones for the binary robot tool segmentation task.
-In this early version of the paper, we present the dataset for the SegSTRONG-C challenge along with its collection and annotation pipeline, the evaluation metrics for the challenge, and a baseline analysis with some segmentation networks and data augmentation methods as baseline models.
-Dataset
+Surgical data science has seen rapid advancement due to the excellent performance of end-to-end deep neural networks (DNNs) for surgical video analysis.
+Despite their successes, end-to-end DNNs have been proven susceptible to even minor corruptions, substantially impairing the model’s performance.
+This vulnerability has become a major concern for the translation of cutting-edge technology, especially for high-stakes decision-making in surgical data science.
+We introduce SegSTRONG-C, a benchmark and challenge in surgical data science dedicated, aiming to better understand model deterioration under unforeseen but plausible non-adversarial corruption and the capabilities of contemporary methods that seek to improve it.
+Built on a dataset generated through counterfactual robotic replay, SegSTRONG-C provides paired clean and corrupted surgical video sequences, enabling reproducible evaluation of model robustness.
+Participants are challenged to train tool segmentation algorithms exclusively on uncorrupted data but are evaluated on corrupted test domains for the task of binary robotic tool segmentation.
+Through comprehensive baseline experiments and participating submissions from widespread community engagement, SegSTRONG-C reveals key themes for model failure and identifies promising directions for improving robustness.
+The performance of challenge winners, achieving an average 0.9394 DSC and 0.9301 NSD across the unreleased test sets with corruption types: bleeding, smoke, and low brightness, shows inspiring improvement of 0.1471 DSC and 0.2584 NSD in average comparing to strongest baseline methods with UNet architecture trained with AutoAugment.
+This highlights how prior knowledge, customized training strategies, and architectural designs can be leveraged to improve robustness without requiring additional data collection.
+In conclusion, the SegSTRONG-C challenge has identified some practical approaches for enhancing model robustness, yet most approaches relied on conventional techniques that have known, and sometimes quite severe, limitations.
+Looking ahead, we advocate for expanding intellectual diversity and creativity in non-adversarial robustness beyond data augmentation or training scale, calling for new paradigms that enhance universal robustness to corruptions and may enable richer applications in surgical data science.
+keywords:
+Non-adversarial robustness, Surgical data science, Robot tool segmentation
+1 Main
+Surgical data science is now a well-established subfield of AI-assisted medicine and has enjoyed increasing popularity and opportunities through the rapid advancement of end-to-end deep neural networks (DNNs) for image and video processing.
+Tool segmentation (robotic or otherwise) in surgical videos is among the most fundamental surgical data science tasks because it enables numerous downstream analyses and applications, such as gesture and phase recognition [1, 2, 3, 4, 5], skill assessment [6, 7], mixed reality feedback [8, 9, 10, 11, 12] and surgical automation [13, 14], among others.
+DNNs [15, 16, 17, 18, 19, 20, 21, 22, 23] have dominated robot tool segmentation ever since the emergence of deep learning, fueled in large parts by the introduction of public and well annotated datasets of surgical scenes together with the associated community grand challenges, many of which under the EndoVis umbrella [24, 25].
+These challenges have established that, generally, methods that achieve state-of-the-art performance on general computer vision benchmarks [26, 27] also exhibit excellent performance on surgical video.
+However, all of these previous results make the strong assumption that training and test data are highly similar and do not contain any (or at least any considerable) corruption to image quality.
+This means that current challenges cannot assess the non-adversarial robustness of contemporary tool segmentation algorithms to various corruptions of image quality, such as tool occlusion through smoke or bleeding.
+This is a severe shortcoming of current benchmarks because in surgical video, these corruptions are not the exception – they are the norm.
+Because it is now well established that DNNs are highly susceptible to even minor corruptions resulting in considerably impaired performance [28, 29, 30], it is critical to better understand this vulnerability and the approaches one may take to enhance model robustness.
+Various strategies [29] have been proposed to attempt to increase robustness under non-adversarial corruption.
+However, due to the lack of evaluation benchmarks, the best practice for developing non-adversarially robust algorithms for surgical video analysis in general and robot tool segmentation in particular remains unestablished.
+The current datasets and data collection/synthesis strategies used to evaluate non-adversarial robustness are insufficient.
+The most common approaches, chief among them ImageNet-C [31], rely on synthetically crafted corruptions based on common augmentation techniques (such as noise injection and contrast variations) [29].
+These methods do not accurately reflect the interactions of various corruption factors during image formation (such as reduced brightness not only affecting intensity but also noise due to higher sensor sensitivity) [32, 33] and do not offer sufficiently realistic simulation of the corruptions most relevant to robotic surgery and surgical data science.
+In this work, we present the SegSTRONG-C benchmark to draw research attention to the important issue of increasing non-adversarial robustness in surgical video analysis and to evaluate the robustness of contemporary and new approaches under non-adversarial corruptions that commonly happen in the surgical scene.
+This approach contributes to a more trustworthy assessment and analysis of where surgical tool segmentation algorithms stand when dealing with potential corruptions of image quality.
+The SegSTRONG-C challenge introduces a dataset that provides robotic surgical scenes with and without corruptions that are created in a counterfactual setting by leveraging the automation capabilities afforded by robotic systems.
+Specifically, we create a dataset comprising paired surgical video sequences under clean and corrupted conditions, that is created by replaying the same robotic actions under a multitude of corruption conditions: 1) Uncorrupted video sequences, 2) variation in background, 3) variation in brightness, 4) presence/absence of blood, and 5) presence/absence of smoke.
+In contrast to all prior work, these corruptions are not synthetically simulated but introduced in the real world, resulting in individual corruptions potentially affecting multiple different image properties (e.g., reduced brightness potentially affecting noise, as in the example above).
+This dataset builds the foundation for the SegSTRONG-C challenge, where participants are challenged to develop algorithms solely on uncorrupted sequences but are evaluated and ranked based on their performance on corrupted videos for the robot tool segmentation task.
+We conducted precisely controlled baseline experiments with various representative DNNs, including “naively trained” convolutional neural networks (CNNs), transformers, and pre-trained foundation models, along with improvement attempts like auxiliary data and general data augmentation.
+We summarize the main findings here and provide a thorough quantitative and qualitative analysis in Section 2.1:
+  • Non-adversarial corruptions lead to substantial performance degradation in end-to-end DNNs.
+  • Foundation models exhibit stronger robustness than DNNs to specific corruptions.
+  • Domain shift and reduced visibility make major contributions to the performance degradation.
+  • Acquiring sufficient data from corrupted domains can mitigate domain shift but poses practical challenges.
+  • General data augmentations enhance feature extraction under non-adversarial corruptions.
+Our challenge received high community participation in a 4-month running window with 40 registrations and 10 final submissions.
+All submissions applied different solutions that can be categorized into three main aspects: customized data augmentation, architecture design, and geometry-aware approaches.
+Here, we summarize the main findings, consolidating our observations across all submissions and detail quantitative and qualitative results in Section 2.2:
+  • Strong robustness can be achieved without access to corruption-domain data when the corruption type is known.
+  • Customized data augmentations are, for now, the most common approach to attempting robustness.
+  • Strong architecture is a necessary but not sufficient condition for achieving robustness.
+  • Geometry-aware approaches represent a promising direction for enhancing robustness.
+In summary, the SegSTRONG-C challenge- a benchmark targeted at a comprehensive evaluation of the non-adversarial robustness of contemporary tool segmentation models- confirmed that non-adversarial robustness is an important concern for vision-based surgical data science.
+The severe performance deterioration of the “naively trained” baseline models emphasizes the need for more targeted development efforts to increase robustness, or at least, dedicated evaluations to assess non-adversarial robustness, especially in more translational research.
+The challenge also identified several strategies that are reasonably practical to enhance model robustness.
+They include model selection and training design, such as relying on foundation models or training models with larger capacity and introducing customized and general data augmentations.
+Augmentations especially offer utility if the corruption type is well-known and can be simulated with reasonable fidelity as part of data augmentation.
+However, breaking the assumption that the corruption type is known during model development presents a more complex challenge, the solution to which will require a more universal level of non-adversarial robustness.
+Approaches to achieve this level of robustness are being attempted, e.g., through the incorporation of geometric constraints or other forms of prior knowledge.
+The geometry-aware approach introduced by one winning team leverages universal low-level visual representations (e.g., depth) to improve model robustness, assuming these representations can be extracted robustly.
+Overall, we argue for increasing the intellectual diversity and creativity of new approaches to address non-adversarial robustness beyond data augmentation or training set scale.
+Ultimately, we believe that these new paradigms will not only provide enhanced and more universal robustness to corruptions of any kind but, through their more complete modeling (e.g., by incorporating robot kinematics or other multi-modal data sources), enable richer downstream applications to benefit surgical data science as a whole.
+2 Results
+2.1 Baseline Analysis
+To establish reference performance and reveal fundamental challenges in robust segmentation under corruption, we conducted a baseline analysis using carefully selected representative models.
+We included UNet[16] and DeepLabv3plus[17] as popular choice examples of CNNs and SETR[20] and Segformer[21] as representatives of transformer-based architectures.
+All models were trained in a fully supervised manner on uncorrupted training sequences in line with the challenge requirements.
+We apply AutoAugment [34] during UNet training to benchmark the utility of general rule-based augmentation methods.
+In addition, we included the Segment Anything Model 2 (SAM2)[35], as reported in [36], as a representative foundation model to evaluate its zero-shot performance on corrupted surgical data.
+This inclusion allows us to explore the potential and current limitations of large-scale pre-trained models in the domain of surgical video segmentation.
+The following sections detail the performance of these baselines across various corruption types and summarize key insights into model degradation patterns and mitigation strategies.
+Table 1: Experiment results of baseline models across different domains.
+DSC Score uncorrupted Background Bleeding Smoke Low
+Change Brightness
+DeepLabv3+ [17] 0.8961 0.8102 0.6896 0.6538 0.5352
+Segformer [21] 0.8993 0.7865 0.6802 0.6906 0.6145
+SETR [20] 0.8381 0.8337 0.5064 0.6657 0.0275
+UNet [16] 0.9410 0.8645 0.7052 0.6603 0.5750
+UNet [16] + AutoAug [34] 0.9568 0.9446 0.7910 0.8895 0.6965
+SAM2 [35] 0.9096 0.9325 0.8628 0.8660 0.4373
+NSD Score uncorrupted Background Bleeding Smoke Low
+Change Brightness
+DeepLabv3+ [17] 0.7941 0.6140 0.5629 0.4637 0.4000
+Segformer [21] 0.8023 0.5962 0.5133 0.5266 0.4194
+SETR [20] 0.6286 0.6276 0.3312 0.4409 0.0092
+UNet [16] 0.8888 0.7020 0.5677 0.5084 0.4390
+UNet [16] + AutoAug [34] 0.9542 0.9196 0.6654 0.8152 0.5344
+SAM2 [35] 0.8479 0.9186 0.8002 0.7808 0.2882
+Baseline performance and failure analysis
+Non-adversarial corruptions lead to substantial performance degradation in end-to-end DNNs.
+As shown in Table 1, all four representative end-to-end DNNs—UNet, DeepLabv3+, SETR, and Segformer—achieve strong performance on the uncorrupted test set, with an average DSC of 0.8936 and NSD of 0.7785.
+However, their performance exhibits a substantial drop when evaluated on corrupted test sequences, including background change, bleeding, low brightness, and smoke.
+Across all models, the average performance is 0.8238 DSC and 0.6350 NSD for the background change domain, 0.6454 DSC and 0.4938 NSD for the bleeding domain, 0.6676 DSC and 0.4849 NSD for the smoke domain, and 0.4385 DSC and 0.3165 NSD for the low brightness domain, respectively.
+This indicates that non-adversarial corruptions substantially affect both segmentation accuracy and boundary precision.
+Foundation models exhibit stronger robustness than DNNs to specific corruptions.
+Foundation models are known for their strong zero-shot generalization.
+When evaluated on the uncorrupted domain, SAM2 [35] achieves performance comparable to supervised DNNs, with 0.9096 DSC and 0.8479 NSD, using only a single point prompt on the first frame of each sequence.
+Notably, SAM2 outperforms the baselines under background change corruption and maintains less significant performance degradation under bleeding and smoke corruptions.
+However, SAM2 is highly susceptible to low brightness, showing 0.4373 DSC and 0.2882 NSD.
+Figure 1: Qualitative comparison among baseline models. The area with green mask denotes true positive. The red area denotes false positives. The blue area denotes false negatives.
+Table 2: Out of distribution study results.
+DSC Score Training uncorrupted Background Bleeding Smoke Low
+Domain Change Brightness
+UNet [16] uncorrupted 0.9410 0.8645 0.7052 0.6603 0.5750
+UNet [16] Test 0.9410 0.8765 0.8651 0.8689 0.7866
+UNet [16] All 0.8843 0.8492 0.8084 0.8718 0.7629
+UNet [16] + AutoAug [34] uncorrupted 0.9568 0.9446 0.7910 0.8895 0.6965
+UNet [16] + AutoAug [34] Test 0.9568 0.9464 0.9332 0.9441 0.8694
+UNet [16] + AutoAug [34] All 0.9590 0.9522 0.9507 0.9565 0.9059
+NSD Score Training uncorrupted Background Bleeding Smoke Low
+Domain Change Brightness
+UNet [16] uncorrupted 0.8888 0.7020 0.5677 0.5084 0.4390
+UNet [16] Test 0.8888 0.7691 0.7243 0.7616 0.6463
+UNet [16] All 0.7442 0.7187 0.6768 0.7343 0.5928
+UNet [16] + AutoAug [34] uncorrupted 0.9542 0.9196 0.6654 0.8152 0.5344
+UNet [16] + AutoAug [34] Test 0.9542 0.9211 0.8888 0.9052 0.7626
+UNet [16] + AutoAug [34] All 0.9581 0.9476 0.9306 0.9524 0.8037
+Domain shift and reduced visibility make major contributions to the performance degradation.
+We conducted a detailed failure analysis of the baseline models to better understand the causes of performance degradation under non-adversarial corruptions.
+We begin with a qualitative assessment, as illustrated in the examples shown in Figure 1.
+Each corruption type introduces distinct failure modes.
+In the background change domain, the appearance of the robot tool remains similar to the uncorrupted domain; however, models often misclassify bright regions in the altered background as surgical tools, resulting in false positives.
+A similar phenomenon is observed under smoke corruption, where translucent white areas introduced by smoke are incorrectly segmented as tool regions.
+In contrast, bleeding occludes parts of the tool, leading to false negatives, while low brightness reduces overall visibility and similarly causes false negative predictions.
+Based on these observations – and noting that baseline models were trained only on uncorrupted data – we hypothesize that performance degradation is driven by two major factors: (i) domain shift introduced by corruption, and (ii) reduced visibility, arising from more complex background, occlusions, reduced contrast, and visual ambiguity.
+To disentangle these effects, we designed an experiment that isolates the role of domain shift.
+Specifically, we retrained the UNet model using two additional training setups: (1) using only samples from the same corruption domain as the test set and (2) using samples from all corruption domains.
+The results are reported in Table 2.
+In the background change domain, models trained on corrupted data (either from the same domain or across all domains) did not outperform the model trained only on uncorrupted data.
+This suggests that performance degradation in this case is primarily attributable to feature extraction difficulty rather than domain shift.
+In contrast, for bleeding, smoke, and low brightness, training on corrupted domains yielded a significant performance improvement—exceeding 0.1 DSC and NSD gains compared to the uncorrupted-trained model.
+However, these models still did not reach the performance levels observed on the uncorrupted test set, indicating that both domain shift and feature extraction challenges contribute substantially to the observed degradation in these corruption types.
+Improvement strategies
+We explore straightforward strategies to improve non-adversarial robustness, guided by the insights from our failure analysis.
+Figure 2: Auxiliary data study results. The x-axis denotes the ratio of the auxiliary data and uncorrupted data. The y-axis denotes the DSC/NSD scores.
+Acquiring sufficient data from corrupted domains can mitigate domain shift but poses practical challenges.
+Incorporating training data from corrupted domains is a natural strategy for mitigating performance degradation caused by domain shift.
+However, since such corruptions are less frequent and harder to capture systematically, data from these domains is typically scarce.
+In this section, we investigate how much auxiliary corrupted data is required to overcome the effects of domain shift associated with each corruption type.
+We divide the training set into five subsets based on camera and robot configurations, as summarized in Table 5.
+Using UNet and UNet + AutoAugment as representative models, we conduct experiments in the bleeding, smoke, and low brightness domains—where domain shift effects are most pronounced.
+For each corruption type, the training data always includes uncorrupted samples from all five subsets, and we incrementally add one subset of corrupted data at a time as auxiliary training data.
+We also include a variant trained with corrupted data from all corrupted domains.
+The results are shown in Figure 2.
+The yellow line represents a model trained only on corrupted-domain data (across all five subsets), serving as an upper bound where domain shift is absent.
+Across all three corruption types, even small amounts of auxiliary corrupted data consistently improved performance.
+With a sufficient amount—roughly over 50% of the uncorrupted training data scale—the model’s performance approached that of the domain-aligned upper bound.
+However, the volume of corrupted data required to achieve this performance highlights a key limitation: Solving domain shift purely through additional data acquisition is feasible in principle but difficult in practice due to the scarcity and variability of the corruptions.
+General data augmentations enhance feature extraction under non-adversarial corruptions.
+AutoAugment[34] applies an ensemble of general image augmentations, including color-based transformations (e.g., contrast adjustment) and spatial manipulations (e.g., rotation).
+In Table 1, the UNet + AutoAugment model demonstrates an improvement over vanilla UNet, with a 0.0158 DSC and 0.0654 NSD gain on the uncorrupted test set.
+More importantly, AutoAugment improves the average performance under corruption by 0.1292 DSC and 0.1794 NSD across all corruption types.
+This effect is particularly evident in the smoke domain, where the performance improved in DSC from 0.6603 to 0.8895 and in NSD from 0.5084 to 0.8152.
+Additionally, UNet with AutoAugment outperforms SAM2 on low brightness, suggesting enhanced resilience in more visually degraded scenarios.
+Further results in Table 2 reinforce this trend: UNet + AutoAugment, when trained on either the corrupted domain or across all domains, consistently outperforms the UNet baselines—by an average of 0.1009 DSC and 0.1981 NSD.
+Notably, since domain shift is not present in these training settings, the performance gains can be attributed to improved feature extraction under challenging visual conditions.
+However, we observe no clear evidence that AutoAugment alone mitigates the effects of domain shift.
+2.2 Challenge Evaluation
+We evaluated all participant submissions on both the uncorrupted test domain and the corrupted test domains, which include background change, bleeding, smoke, and low brightness.
+Challenge rankings were determined based solely on performance on the latter three domains, as they remained unreleased during the competition to ensure fairness.
+Based on the cumulative ranking points shown in Table 3, Team ICVS-2AI achieved the highest score with 58 points, securing first place.
+Team UBCRCL followed with 55 points, placing second, and Team Talor ranked third with 43 points.
+The following paragraphs summarize the key findings derived from the evaluation results and the corresponding approaches summarized in Section 5.5.
+Table 3: Ranking of all participants.
+Teams NSD DSC Total Final
+Bleeding Smoke LB Bleeding Smoke LB Points Rank
+ICVS-2AI 9 10 10 9 10 10 58 1
+UBCRCL 10 8 9 10 9 9 55 2
+Tailor 7 7 8 7 7 7 43 3
+FightTumor 6 9 6 5 8 6 40 4
+AIMI 6 5 7 6 5 8 37 5
+SK 8 6 2 8 6 2 32 6
+SAM_KHU 4 3 5 4 3 4 23 7
+CCG_DGIST 3 4 3 3 4 4 21 8
+VSI 2 2 4 2 2 5 17 9
+Medical_Mechatronics 1 1 1 1 1 1 6 10
+Table 4: Test results of all participants.
+DSC Score uncorrupted BG Change Bleeding Smoke Low Brightness
+SAM2 [35] 0.8479 0.9186 0.8002 0.7808 0.2882
+UNet [16] + AutoAugment [34] 0.9568 0.9446 0.7910 0.8895 0.6965
+VSI 0.6547 0.6161 0.6021 0.6153 0.5805
+UBCRCL 0.9587 0.9564 0.9424 0.9426 0.9030
+SAM_KHU 0.8476 0.8023 0.7644 0.7718 0.5871
+Medical_Mechatronics 0.6607 0.7625 0.5079 0.5670 0.4372
+CCG_DGIST 0.9215 0.9008 0.7335 0.8474 0.5796
+SK 0.9402 0.9493 0.9006 0.9062 0.5365
+Tailor 0.9439 0.9432 0.8902 0.9329 0.8211
+FightTumor 0.9569 0.9486 0.8528 0.9453 0.7465
+ICVS-2AI 0.9706 0.9530 0.9151 0.9657 0.9102
+AIMI 0.9626 0.9508 0.8517 0.8625 0.7895
+NSD Score uncorrupted BG Change Bleeding Smoke Low Brightness
+SAM2 [35] 0.8479 0.9186 0.8002 0.7808 0.2882
+UNet [16] + AutoAugment [34] 0.9542 0.9196 0.6654 0.8152 0.5344
+VSI 0.5984 0.5411 0.5023 0.5623 0.4850
+UBCRCL 0.9699 0.9690 0.9345 0.9340 0.8268
+SAM_KHU 0.7531 0.6759 0.6500 0.6540 0.4277
+Medical_Mechatronics 0.3537 0.4969 0.2627 0.3816 0.2137
+CCG_DGIST 0.8523 0.7872 0.5816 0.6882 0.4312
+SK 0.8937 0.9244 0.8099 0.8111 0.3173
+Tailor 0.9178 0.9176 0.7884 0.8874 0.6353
+FightTumor 0.9598 0.9279 0.7636 0.9240 0.6116
+ICVS-2AI 0.9934 0.9472 0.8870 0.9827 0.8730
+AIMI 0.9693 0.9192 0.7692 0.7763 0.6437
+Strong robustness can be achieved without access to corruption-domain data when the corruption type is known.
+Despite not having access to training data from the corrupted domains, the top-performing teams achieved robustness comparable to that of the UNet + AutoAugment model trained on data from all domains.
+In the background change and bleeding domains, Team UBCRCL achieved the best results, with DSC scores of 0.9564 and 0.9424 and NSD scores of 0.9690 and 0.9345, respectively.
+In the smoke and low brightness domains, Team ICVS-2AI led with DSC scores of 0.9657 and 0.9102 and NSD scores of 0.9827 and 0.8730, respectively, showing substantially improved robustness compared to the baseline models, such as SAM2 and UNet + AutoAugment.
+Notably, both winning teams reported customizing their approaches based on the pre-known corruption types, highlighting the value of targeted, corruption-aware model design in achieving robust performance without requiring domain-aligned training data.
+Customized data augmentations are, for now, the most common approach to attempt robustness.
+Most teams leveraged prior knowledge of the corruption types to implement corruption-specific augmentation strategies.
+For the smoke corruption, team UBCRCL, SK, Tailor, FightTumor, and ICVS-2AI apply various specialized smoke augmentations, and all achieve an average of 0.9385 DSC score and 0.9078 NSD score.
+On the contrary, without specialized smoke augmentation, team CCG_DGIST and team AIMI, although achieving comparable performance on the uncorrupted domain, only show an average of 0.8549 DSC score and 0.7323 NSD score.
+Unlike others that generate fake smoke in rule-based ways, team ICVS-2AI trains a CycleGAN from internal real data for the smoke generation, resulting in the best performance (0.9657 DSC score and 0.9827 NSD score) in the smoke domain.
+For bleeding corruption, Team SK used red-colored ellipses to simulate blood, while Team UBCRCL applied general occlusion-based augmentations with a similar masking effect.
+These strategies yielded the relative higher performance of 0.9215 DSC and 0.8722 NSD on average.
+Other teams with comparable uncorrupted-domain performance (Tailor, ICVS-2AI, FightTumor, CCG_DGIST, and AIMI) did not incorporate bleeding-specific augmentation and experienced significantly larger performance deterioration, achieving only 0.8487 DSC and 0.7580 NSD on average.
+In the low-brightness domain, the top three performers—UBCRCL, ICVS-2AI, and Tailor—also had the smallest performance degradation.
+UBCRCL and Tailor explicitly applied brightness reduction during training, while ICVS-2AI did not use low-brightness-specific augmentation but still performed well, likely due to overall strong model capacity.
+Team SK attempted a model-based de-artifacting approach, but it proved less effective in this domain.
+Even among the top teams, performance deterioration under low brightness remained non-negligible, suggesting that visibility degradation imposes limitations that augmentation alone cannot fully overcome.
+Figure 3: Correlations between uncorrupted test performance and corrupted test performance. Numbers in the legend show the quantitative correlation.
+Strong architecture is a necessary but not sufficient condition for robustness.
+Since all participants trained their models solely on the uncorrupted domain, we use performance on the uncorrupted test set as a proxy for the underlying architectural capacity.
+As shown in Figure 3, there is a strong positive correlation between performance on the uncorrupted domain and on most corrupted test domains (background change, bleeding, and smoke) for both DSC and NSD scores.
+Given that models were not exposed to corrupted training data or extensive augmentation, this correlation suggests that architectures with higher capacity are more likely to retain performance under certain corruption types.
+In the low brightness domain, two notable outliers—Team SK and CCG_DGIST—achieved strong, uncorrupted domain performance but experienced substantial degradation under corruption.
+Neither team reported using general data augmentation during training.
+This underscores that while a strong architecture provides a solid foundation, it is not sufficient for robustness.
+Geometry-aware approaches represent a promising direction for enhancing robustness.
+Team UBCRCL explored the incorporation of geometric information (i.e., depth) by leveraging pseudo-depth maps inferred from off-the-shelf foundation models.
+Their method encodes both RGB and depth representations into a shared latent feature space for final prediction, enabling a form of RGBD-based segmentation.
+This approach demonstrated strong robustness across corruptions.
+Despite using a less powerful architecture compared to Team ICVS-2AI—with a 0.0119 lower DSC and 0.0235 lower NSD score on the uncorrupted domain—Team UBCRCL achieved the best performance under background change corruption, with a DSC of 0.9564 and NSD of 0.9690, comaprable to the performance on uncorrupted domains (0.9587 DSC and 0.9699 NSD).
+These results suggest that incorporating low-level geometric information can help models learn more informative and generalizable features, potentially reducing reliance on explicit corruption-specific augmentation.
+Geometry-aware strategies may, therefore, offer a complementary and robust pathway for addressing domain shift in surgical scene understanding.
+3 Discussion
+While benchmark performance remains a standard metric for algorithm development, it alone is insufficient to evaluate model reliability under real-world surgical conditions.
+As our baseline analysis confirmed, “naively trained” end-to-end DNNs that perform well on the uncorrupted domain often suffer substantial degradation when exposed to non-adversarial corruptions such as smoke, bleeding, low lighting, and background variation.
+In the context of high-stakes surgical applications, overlooking this robustness concern during development can result in serious clinical consequences.
+Our work addresses this gap by introducing SegSTRONG-C, a benchmark and challenge focused specifically on robustness in robotic tool segmentation.
+In addition to evaluating existing methods, we propose a reproducible data generation pipeline using robotic trajectory replay to systematically simulate corrupted test data—laying the foundation for future benchmarks targeting diverse clinical scenarios.
+To better understand the nature of performance degradation, we disentangled two contributing factors: domain shift and reduced visibility.
+Through customized training on corrupted data, we confirmed that domain shift plays a major role in performance deterioration.
+However, even with domain-aligned training, performance gaps persist, highlighting the importance of enhancing models’ ability to extract meaningful features from visually degraded inputs.
+Our baseline analysis revealed that while incorporating corrupted data can fully mitigate domain shift, doing so requires large-scale collection efforts—often exceeding 50% of the uncorrupted training set—which may not be feasible in most clinical datasets.
+As a more practical alternative, general data augmentation, such as AutoAugment, offers a cost-effective and annotation-free approach to improve feature extraction.
+While no evidence shows that these methods resolve domain shift, they enhance the model’s ability to extract features under reduced visibility to further improve their robustness to a limited degree under corruptions like smoke and low contrast.
+Through the SegSTRONG-C challenge, we observed that top-performing teams achieved robust performance without access to corrupted-domain data, primarily by leveraging prior knowledge of the corruption types.
+Customized data augmentations emerged as the most common and effective strategy, enabling targeted mitigation of domain shift.
+For example, generating synthetic smoke using CycleGAN or simulating bleeding through occlusion-based augmentations yielded significant gains.
+These approaches provide a compelling middle ground: they require minimal real corrupted data but achieve robustness close to fully supervised models.
+Beyond augmentation, our results also underscore the importance of architecture choice.
+Models with higher performance on the uncorrupted domain generally achieved stronger results on corrupted domains.
+This correlation highlights the role of strong feature extraction backbones.
+Despite these achievement, the SegSTRONG-C challenge leaves several open questions.
+All corruption types were disclosed in advance, allowing teams to engineer targeted solutions.
+However, in real surgical environments, the nature and severity of visual corruptions are not always known.
+As such, the robustness of current submissions to unknown or unseen non-adversarial corruptions remains less explored.
+Team UBCRCL, however, not only applied customized approaches but also attempted a geometry-aware method.
+They incorporated pseudo-depth generated from foundation models (e.g., DepthAnything) to form RGB-D feature encodings, yielding top-tier performance on several corruption types.
+This suggests that geometric representations provide a more invariant, unified view of the scene, which generalizes more effectively under corruption.
+The rise of foundation models capable of extracting such geometric priors with strong zero-shot generalization makes this direction not only effective but also practical for real-world use.
+Other ongoing research efforts aiming for more generalizable feature extraction, such as domain generalization [37, 38, 39], also merit validation.
+Novel paradigms, such as multi-modal alignment [40, 41] and digital twin [4, 2, 13, 42, 40] may also provide opportunities for more universal and comprehensive representation that enable systematic advancement of research and clinical translation in surgical data science.
+4 Conclusion
+We presented SegSTRONG-C, a benchmark and community challenge in surgical data science focused on evaluating algorithmic robustness under realistic, non-adversarial corruptions.
+By simulating realistic visual challenges encountered during surgery, SegSTRONG-C fills a critical gap in the field and provides an essential tool for evaluating deep learning systems in corrupted conditions that closely mirror real-world surgical environments.
+The results from SegSTRONG-C provide a comprehensive assessment of robustness strategies in surgical data science.
+Our findings demonstrate that strong baseline models degrade significantly under corruption and that robustness can be substantially improved- without access to corrupted data- through customized augmentation, geometry-aware designs, and architecture selection.
+These insights offer clear, actionable guidance for building future surgical AI systems with strong performance and robustness.
+By identifying effective practices and exposing the limitations of current approaches, SegSTRONG-C paves the path toward clinically dependable and corruption-aware surgical intelligence.
+5 Method
+5.1 Dataset
 The dataset consists of mock endoscopic video sequences with corresponding binary segmentation masks for the robot tool segmentation task.
-We mock the endoscopic scene with two patient-side manipulators (PSMs) and animal tissue backgrounds to ensure a photo-realistic appearance.
+We mock the endoscopic scene with two patient-side manipulators (PSMs) from the da Vinci surgical robot and animal tissue backgrounds to ensure a photo-realistic appearance.
 We manually teleoperate the robot to generate the trajectory in free space.
@@
 We use the final 3 sequences with certain corruptions(smoke, bleeding, and low brightness) as the test set.
-The models submitted by the challenge participants will be tested on this test set consisting of sequences with photo-realistic non-adversarial corruptions.
-The models will be tested on each corruption separately.
-As the challenge is currently active, we withhold the release of the test set.
+The models submitted by the challenge participants are tested on this test set consisting of sequences with photo-realistic non-adversarial corruptions.
+The models are tested on each corruption separately.
 Table 5 shows the summary of the dataset.
-All corrupted versions for the training and validation will be released after the challenge submission.
+All corrupted versions for the training and validation are released and can be found on the challenge website segstrongc.cs.jhu.edu.
 The example images are shown in Figure 4.
@@
 1 14 300 + 300 300 + 300 - - -
-Test set 9 15 - - 300 + 300 300 + 300 300 + 300
-9 16 - - 300 + 300 300 + 300 300 + 300
-9 17 - - 300 + 300 300 + 300 300 + 300
-Figure 1: Example images for regular image and non-adversarial corruptions(background, smoke, bleeding, and low brightness.).
+Test set 9 15 300 + 300 300 + 300 300 + 300 300 + 300 300 + 300
+9 16 300 + 300 300 + 300 300 + 300 300 + 300 300 + 300
+9 17 300 + 300 300 + 300 300 + 300 300 + 300 300 + 300
+Figure 4: Example images for uncorrupted image and non-adversarial corruptions(background, smoke, bleeding, and low brightness.).
 Data Collection
-The data was acquired in the Robotorium of the Laboratory for Computational Sensing and Robotics (LCSR), Johns Hopkins University by surgical robotics experts familiar with operating the da Vinci robot via dVRK [18].
+The data was acquired in the Robotorium of the Laboratory for Computational Sensing and Robotics (LCSR), Johns Hopkins University, by surgical robotics experts familiar with operating the da Vinci robot via dVRK [43].
 We use dVRK [43] as our robot platform with the endoscopic camera manufactured by SCHÖLLY and the image process unit manufactured by Ikegami as the perception part.
+We refer to the data collection pipeline in CaRTS [40, 41] to generate paired corrupted video sequences.
 We first adjust the camera/robot configuration and perform calibrations.
 For each configuration, we collect different trajectories by human operation.
-For each trajectory, we replay the recorded kinematics to reproduce the trajectory and record video sequences under different scenarios including non-corrupted and corrupted versions with background, smoke, bleeding, and low brightness.
+For each trajectory, we replay the recorded kinematics to reproduce the trajectory and record video sequences under different scenarios, including non-corrupted and corrupted versions with background, smoke, bleeding, and low brightness.
 To generate non-adversarial corruptions for the same testing samples, we record the robot trajectory, replay the kinematics via dVRK, and manually add corruption for each replay.
@@
 Data annotation
-The annotations are generated via a semi-automatic pipeline based on exclusively collected sequences with purely black backgrounds where the only salient areas in the images are the PSMs.
-We first automatically generate a segmentation mask for the PSMs via the segment anything model (SAM) [36] where the prompts are generated via traditional background extraction algorithm.
-We select experts who are familiar with the robot and segmentation task to be the annotators to verify and correct the automatically generated annotation.
+The annotations are generated via a semi-automatic pipeline based on exclusively collected sequences with purely black backgrounds, where the only salient areas in the images are the PSMs.
+We first automatically generate a segmentation mask for the PSMs via the segment anything model (SAM) [44], where the prompts are generated via a traditional background extraction algorithm [45].
+We select experts who are familiar with the da Vinci robot and segmentation task to be the annotators to verify and correct the automatically generated annotation.
 The annotation pipeline can be expressed in the following steps:
-  • Step 1: Prompt generation - We use a traditional background extraction algorithm to generate rough masks for the foreground PSMs.
-The rough masks are converted to the prompt points and bounding boxes for SAM.
+  • Step 1: Prompt generation - We use a traditional background extraction algorithm to generate coarse masks for the foreground PSMs.
+The coarse masks are converted to the prompt points and bounding boxes for SAM.
   • Step 2: Automatic generation - We input the prompt with the image to generate a fine mask for the robot tool. The first two steps are fully automatic.
-  • Step 3: Failure case selection - Human annotators are involved in examining the generated mask from SAM and selecting the failure cases.
+  • Step 3: Failure case selection - Human annotators are involved in examining the mask generated from SAM and selecting the failure cases.
 We had two annotators examining the same sequences and using the union of their selections as the set of failure cases.
@@
 If the annotator had three attempts but did not get satisfactory results, they would draw the contour to annotate this sample manually.
-Evaluation
+5.2 Evaluation
 We introduce how we quantitatively evaluate the performance of an algorithm and how we decide the ranking of multiple algorithms that participate in the challenge.
-Metrics:
-We use dice similarity coefficient (DSC) and normalized surface distance (NSD) averaged from different tolerances for the robot tool for multiple corruptions - low brightness, smoke, and blood.
+Metrics
+We use the DICE similarity coefficient (DSC) and normalized surface distance (NSD) averaged from different tolerances for the robot tool for multiple corruptions - low brightness, smoke, and blood.
 The DSC is a widely used metric in the field of medical image analysis.
@@
 $DSC=\frac{2TP}{2TP+FP+FN}$
-, where $TP$ is the number of true positive pixels for the class, $FP$ is the number of false positives and $FN$ is the number of false negatives.
+, where $TP$ is the number of true positive pixels for the class, $FP$ is the number of false positives, and $FN$ is the number of false negatives.
 NSD measures the overlap of two boundaries between the predicted and ground truth segmentation masks.
-A boundary pixel is counted as overlapping when the closest distance to the other boundary is less or equal to the specified tolerance.
+A boundary pixel is counted as overlapping when the closest distance to the other boundary is less than or equal to the specified tolerance.
 NSD for a 2D binary class image segmentation task is defined as:
@@
 We calculate the average DSC and NSD over images.
-DSC is the standard metric for segmentation while NSD is complementary to DSC.
+DSC is the standard metric for segmentation, while NSD is complementary to DSC.
 DSC reveals performance more on the chunk area while NDS focuses on the boundary.
 We evaluate the performance only on the unseen domain to encourage participants to focus on the algorithm’s robustness.
-Ranking:
+Ranking
 We rank algorithms with total scores summed over 3 test domains - bleeding, smoke, and low brightness.
@@
 Then, we do a pairwise significance test for adjacent methods to test whether they have statistically significant differences in performance.
-If yes, we rank them according to the order, otherwise they have the same rank.
+If yes, we rank them according to the order, otherwise, they have the same rank.
 The mean value indicates the overall performance, and the pairwise significance test can justify whether one method has a significant advantage over the other one.
@@
 Since the goal for the participating team is to develop methods that are robust against all corruptions instead of algorithms achieving dominating performance on one domain but failing to generalize to others, the proposed overall ranking method better suits the goal of the challenge.
-Baseline Analysis
-In this section, we first briefly introduce the selected segmentation networks and data augmentation methods.
-Then, we present the testing results of these methods and perform a brief analysis of the baseline results.
-Networks
-U-Net
-: U-Net[2] is originally developed for biomedical image segmentation.
-The architecture of U-Net consists of two paths: a contracting path and an expansive path.
-The contracting path is characterized by repeated convolutions and resembles a conventional convolutional network, followed by a rectified linear unit(RELU) and a max pooling operation.
-On the other hand, the expansive path enhances feature and spatial information through a series of up-convolutions and concatenation of features from the contracting path.
-Our implementation uses a 4-level vanilla U-Net as the baseline model.
+5.3 Challenge Logistics
+SegSTRONG-C challenge launches as a one-time sub-challenge under the EndoVis2024 challenge [^1] https://opencas.dkfz.de/endovis/challenges/2024/.
+The challenge launches on May 1st, 2024, with all training and validation datasets released.
+All submissions were collected before September 6th, 2024, and evaluated using docker containers.
+We announced one winner and two runner-ups on MICCAI2024 (October 10th, 2024).
+The final results were released on October 12th, 2024, and test data were released on October 18th, 2024.
+Registration and Submissions
+In total, we received 40 registrations from all over the world.
+All related countries are listed below: Canada (12.5%), USA (7.5%), Republic of Korea (10%), China (32.5%), Switzerland(2.5%), Japan(2.5%), Portugal(2.5%), India(10%), Germany(7.5%), UK(2.5%), Iran(2.5%), Denmark(2.5%), Singapore(2.5%), Morocco(2.5%).
+Within all registrations, we received 10 valid submissions.
+They are from Canada (20%), USA (20%), Republic of Korea (20%), China (10%), Switzerland(10%), Japan(10%), Portugal(10%).
+5.4 Baseline Models
+In this section, we introduce the selected segmentation networks and data augmentation methods for baseline analysis.
 DeepLabV3+
-: The most recent version of the DeepLab family is DeepLabV3+ [3].
+The most recent version of the DeepLab family is DeepLabV3+ [17].
 The architecture is changed to an encoder-decoder structure.
@@
 SegFormer
-: SegFormer [7] is a semantic segmentation framework that combines a hierarchical transformer encoder with a multi-layer perceptron (MLP) decoder.
+SegFormer [21] is a semantic segmentation framework that combines a hierarchical transformer encoder with a multi-layer perceptron (MLP) decoder.
 The hierarchically structured transformer encoder avoids the interpolation of positional codes and outputs multi-scaled features.
@@
 SETR
-: SETR [6] provides an alternative perspective to the encoder-decoder-based FCN architecture.
+SETR [20] provides an alternative perspective to the encoder-decoder-based FCN architecture.
 It treats semantic segmentation as a sequence-to-sequence prediction task in which the author splits the image into patches, linearly projects each patch, adds positional embeddings, and feeds the sequence into the transformer encoder.
 Three different decoders are designed to perform pixel-level segmentation, and we implement all three: (1) Naive upsampling which projects the transformer feature to the dimension of class number and directly upsample to the desired image resolution, (2) Progressive upsampling which alternates convolution and upsampling operations instead of one-step upscaling, and (3) Multi-Level feature aggregation which takes feature representations from $M$ transformer layers uniformly, reshapes and aggregates the feature map top-down, fuses features from all channels and upsamples to the full resolution.
-Augmentation
-Figure 2: Example image augmentations
-AutoAugment
-: AutoAugment [30] searches for the best sequence of transformations that improves the performance of a model on a given dataset.
+We apply naive upsampling in our baseline analysis.
+UNet
+UNet[16] was originally developed for biomedical image segmentation.
+The architecture of UNet consists of two paths: a contracting path and an expansive path.
+The contracting path is characterized by repeated convolutions and resembles a conventional CNN, followed by a rectified linear unit(RELU) and a max pooling operation.
+On the other hand, the expansive path enhances feature and spatial information through a series of up-convolutions and concatenation of features from the contracting path.
+Our implementation uses a 4-level vanilla UNet as the baseline model.
+UNet + AutoAugment
+AutoAugment [34] searches for the best sequence of transformations that improves the performance of a model on a given dataset.
 We modify the Pytorch implementation of the AutoAugmentPolicy class and choose the policy learned on ImageNet.
-Elastic
-: Elastic transformation generates displacement vectors for each image pixel based on random offsets.
-It will transform the image given $\alpha$ and $\beta$, where $\alpha$ controls the strength and $\beta$ controls the smoothness of displacements.
-We modify the Pytorch implementation of the ElasticTransform class.
-Let $\alpha_{x}$ and $\alpha_{y}$ be the magnitude of displacements, $\sigma_{x}$ and $\sigma_{y}$ be the smoothness of displacements along $x$ and $y$ axis, and $h$ and $w$ be the height and width of the image.
-Kernel size along $x$ axis is: $k_{x}=\text{int}(8\sigma_{x}+1)$ thus half of the kernel size is $k_{x\text{\_half}}=\frac{1}{2}(k_{x}-1)$, and the displacement vector $d_{x}$ is first initialized as an array of random numbers between $-1$ and $1$.
-Each pixel is defined as the Gaussian blur of $d_{x}$ times $\alpha_{x}$ and divided by image height $h$.
-The 2D Gaussian kernel $h_{x}$ is defined as $h_{x}[m,n]=h_{1}[m]h_{1}[n]^{T}$ where $h_{1}$ is the normalized discrete 1D Gaussian kernel of length $k_{x}$.
-$d_{x}$, $d_{y}$ is defined as $d_{x}[m,n]=\frac{\alpha_{x}}{h}d_{x}[m,n]*h_{x}[m,n]$, $d_{y}[m,n]=\frac{\alpha_{y}}{w}d_{y}[m,n]*h_{y}[m,n]$.
-The complete displacement vector field is obtained by concatenating $d_{x}$ and $d_{y}$.
-The detailed derivations of $h_{1}$, $d_{x}$, and $d_{y}$ are in the equation:
-$h_{1}[m]=\frac{\exp[(-\frac{1}{2}(\frac{-k_{x\text{\_half}}+m}{\sigma_{x}})^{2})]}{\sum_{n=-k_{x\text{\_half}}}^{k_{x\text{\_half}}}\exp[(-\frac{1}{2}(\frac{-k_{x\text{\_half}}+n}{\sigma_{x}})^{2})]}$
-$d_{x}[m,n]=\frac{\alpha_{x}}{h}\sum_{j=-\infty}^{\infty}\sum_{i=-\infty}^{\infty}d_{x}[i,j]\cdot h_{x}[m-i,n-j]$
-$d_{y}[m,n]=\frac{\alpha_{y}}{w}\sum_{j=-\infty}^{\infty}\sum_{i=-\infty}^{\infty}d_{y}[i,j]\cdot h_{y}[m-i,n-j]$
-Projective
-: Projective transformation transforms the image such that it is viewed from a different viewing perspective with a given probability.
-We modify the Pytorch implementation of the RandomPerspective class.
-The four new coordinates of the corners of the image are first randomly generated according to the input distortion scale.
-In perspective transform, each pixel $(x,y)$ in the original image is mapped to its new coordinate by the formula: $(u_{i},v_{i})=(\frac{ax_{i}+by_{i}+c}{gx_{i}+hy_{i}+1},\frac{dx_{i}+ey_{i}+f}{gx_{i}+hy_{i}+1})$ for $i=0,1,2,3$.
-In matrix form, it can be written as:
-$\begin{bmatrix}x_{0}&y_{0}&1&0&0&0&-x_{0}u_{0}-y_{0}u_{0}\\ x_{1}&y_{1}&1&0&0&0&-x_{1}u_{1}-y_{1}u_{1}\\ x_{2}&y_{2}&1&0&0&0&-x_{2}u_{2}-y_{2}u_{2}\\ x_{3}&y_{3}&1&0&0&0&-x_{3}u_{3}-y_{3}u_{3}\\ 0&0&0&x_{0}&y_{0}&1&-x_{0}v_{0}-y_{0}v_{0}\\ 0&0&0&x_{1}&y_{1}&1&-x_{1}v_{1}-y_{1}v_{1}\\ 0&0&0&x_{2}&y_{2}&1&-x_{2}v_{2}-y_{2}v_{2}\\ 0&0&0&x_{3}&y_{3}&1&-x_{3}v_{3}-y_{3}v_{3}\end{bmatrix}\cdot\begin{bmatrix}a\\ b\\ c\\ d\\ e\\ f\\ g\\ h\end{bmatrix}=\begin{bmatrix}u_{0}\\ u_{1}\\ u_{2}\\ u_{3}\\ v_{0}\\ v_{1}\\ v_{2}\\ v_{3}\end{bmatrix}$
-Solve this system to obtain the least-square solution for the coefficients, and map each pixel to its new position by the perspective transformation formula.
-Results
-We evaluate the performance of baseline models with and without data augmentation across different domains.
-While all models show reasonable accuracy under regular conditions, their effectiveness substantially deteriorates when tested in shifted domains such as in the presence of bleeding, low brightness, and smoke.
-The data augmentations applied, including AutoAugment, elastic, and projective transformations, generally failed to improve model performance in these domains.
-Notably, UNet combined with AutoAugment showed a marginal improvement under smoke conditions but this improvement is inconsistent across all domains.
-As shown in Figure 3, under regular conditions, the models adequately segmented tools from the background, but struggled in areas with appearance change and subtle contrast changes.
-Failures are predominantly under bleeding conditions due to the appearance (color) change of the shaft, leading to incorrect segmentation of the shaft.
-Under low brightness conditions, the darker the region is, the more likely a failure in segmentation will occur.
-The pervasive false positives under smoke conditions indicate that the models interpret the bright regions (smoke) as the foreground.
-These results suggest that all baseline models are highly dependent on adequate targets’ saliency to successfully distinguish between the tools and the background.
-They are also sensitive to any appearance change for the tool (bleeding) and background (smoke).
-Table 2: Experiment results of baseline models across different domains.
-Models NSD DSC
-Bleeding Smoke Low Brightness Bleeding Smoke Low Brightness
-DeepLabv3 [3] 0.5629 0.4637 0.4000 0.6896 0.6538 0.5352
-Segformer [7] 0.5133 0.5266 0.4194 0.6802 0.6906 0.6145
-SETR_PUP [6] 0.2531 0.3354 0.2599 0.4717 0.5848 0.4053
-SETR_MLA [6] 0.2798 0.3374 0.1571 0.5358 0.6206 0.3385
-SETR_Naive [6] 0.3312 0.4409 0.0092 0.5064 0.6657 0.0275
-UNet [2] 0.5677 0.5084 0.4390 0.7052 0.6603 0.5750
-UNet [2] + AutoAugment [30] 0.6654 0.8152 0.5344 0.7910 0.8895 0.6965
-UNet [2] + Elastic 0.5622 0.5207 0.3931 0.6910 0.6583 0.5190
-UNet [2] + Projective 0.5661 0.5702 0.4265 0.6903 0.6978 0.5792
-SETR [6] + AutoAugment [30] 0.1172 0.0910 0.0773 0.1918 0.2157 0.0934
-SETR [6] + Elastic 0.2836 0.3363 0.1654 0.5350 0.6192 0.3472
-SETR [6]+ Projective 0.2240 0.3055 0.0552 0.4168 0.5766 0.1447
-Conclusion
-We present the SegSTRONG-C challenge, aiming to assess algorithms’ robustness to unforeseen yet plausible complications in surgical scenarios.
-To generate non-adversarial corruptions we use and refine the data collection and annotation pipeline from CaRTS [16] and collect 17 sequences for training and evaluation.
-We use DSC and NSD as quantitative metrics to evaluate algorithms’ performance on the testing set and propose a balanced ranking method to alleviate the influence of extreme performance on a single domain while emphasizing the statistical significance during ranking.
-The baseline analysis provides baseline performance for participants to refer to while providing a basic insight that the models’ success relies on adequate saliency and that all models are sensitive to appearance change.
-Acknowledgement:
+We apply AutoAugment during the training of UNet for this baseline.
+SAM2
+SAM2 [35] extends the Segment Anything Model (SAM) [44] for video segmentation by incorporating temporal information.
+It retains SAM’s core components—image encoder, prompt encoder, and mask decoder—while introducing a memory encoder and memory bank to capture frame-to-frame consistency.
+The memory mechanism allows SAM2 to leverage past predictions, enhancing temporal coherence in segmentation.
+We adopt SAM2 as a baseline for its strong performance in point prompt-based video segmentation and its ability to model temporal context effectively.
+We use the ground truth mask to sample 5 positive points and 5 negative points for the first frame of each sequence.
+5.5 Challenge Submissions
+Figure 5: Overview of our proposed corruption simulation + segmentation framework. (Team VSI)
+VSI (Yicheng Leng, Seok Bong Yoo, Eung-Joo Lee):
+Team VSI’s framework is illustrated in Figure 5.
+They use PSPNet50  [46] for the surgical tool segmentation.
+To improve the model’s robustness against different types of corruption, they apply image augmentation techniques from ImageNet-C [31] to simulate the smoke and low-brightness cases.
+They ignore the augmentation of the bleeding case as they think the color and shape of red artifacts added onto the image are difficult to determine, and the bleeding area brings minimal pixel-wise impact on both the surgical tool and tissue background.
+UBCRCL (Zijian Wu, Yuxin Chen, Septimiu E. Salcudean):
+RGB-D semantic segmentation methods have been demonstrated to improve the segmentation performance by fusing depth information [47].
+Team UBCRCL tries to explore the performance of an RGB-D method on surgical datasets, especially when confronting corrupted images.
+They adopt a state-of-the-art method, DFormer [48], a pretraining framework to encode RGB-D representations.
+They apply an off-the-shelf DepthAnything [49] model to generate pseudo depth for each frame, as the depths from stereo-matching [50] are not satisfying without rectified images.
+To enhance the model’s robustness, they perform the 3D Common Corruption (3DCC) [51] for data augmentation.
+Specifically, they tune the following augmentation to different degrees during the training: reducing the image brightness, the color and density of fog, and random occlusions.
+SAM_KHU(Samra Irshad, Shadi Albarqouni, Seong Tae Kim):
+Figure 6: (a). VM-UNet architecture. (b) Details of components in VSS block (Team SAM_KHU)
+State space models (SSMs) [52, 53] have recently gained attention for their ability to model long-distance dependencies with linear complexity, improving sequence modeling and computational efficiency, and are now being integrated with architectures like CNNs for tasks such as medical image segmentation [54, 55].
+Team SAM_KHU applies Vision Mamba-UNet (VM-UNet), a recently proposed SSM-based model for this challenge.
+The architecture is shown in Figure 6.
+The encoder includes VSS blocks from VMamba [56] for feature extraction and patch merging operations for downsampling.
+The decoder employs VSS blocks and patch-expanding operations to recover the original size of the segmentation output.
+The skip connection module uses a simple additive operation.
+To make the model robust to procedural artifacts, they use StarGanv2 [57], an image-to-image translation model that learns the mapping between source and target domains to synthesize these patterns.
+Medical_Mechatronics(Yueyi Sun, An Wang, Long Bai, Hongliang Ren):
+Figure 7: General workflow of our methodology. (Team Medical_Mechatronics)
+For Team Medical_Mechatronics, their primary objective is to introduce comprehensive perturbations onto the pristine images, effectively synthesizing a new, realistic dataset with characteristics mimicking smoke, bleeding, and low-light conditions.
+By training on this augmented dataset, they aim to develop a more robust model that can generalize well to unseen test data.
+They have created different sets of datasets for each corruption and trained type-specific models.
+For smoke, they leverage the widely used fog synthesis approach from the robustness benchmark [31] to generate corruption at different severity levels.
+For blood, they employ the implementation from the semi-synthetic approach [58], utilizing the BloodDroplet class to introduce blood droplets onto images.
+For low brightness, they utilize Photoshop’s batch processing capabilities to enhance images and achieve a similar effect.
+They leverage the great generalization ability of pre-trained visual foundation models in their design.
+Specifically, they employ Yolo-World [59] to generate bounding box prompts for Segment Anything Model 2 (SAM 2) [35], which will be fine-tuned with their augmented datasets.
+The overall pipeline of their method is demonstrated in Figure 7.
+CCG_DGIST(Ihsan Ullah, Ho-Gun Ha, Attaullah Khan, Hyunki Lee):
+Figure 8: The overall architecture of the proposed surgical tool segmentation model. (Team CCG_DGIST)
+Team CCG_DGIST utilizes a multi-attention and lightweight model that achieves competitive performance in surgical tool segmentation.
+The overall framework is shown in Figure 8.
+This framework incorporates several attention modules to enhance performance: (1) The Dilated Gated Attention Block (DGA), which enables the model to focus more precisely on the target regions by capturing both global and local information.
+(2) The Inverted External Attention Block (IEA)[60, 61], an efficient external attention mechanism that strengthens the information associations between samples and extracts overarching characteristics from the entire dataset.
+(3) To capture multistage and multiscale information, they introduce attention bridge modules for both channel and spatial dimensions, termed the Channel Attention Bridge Block (CAB) and Spatial Attention Bridge Block (SAB)[62], which generate their respective attention maps.
+Additionally, depthwise separable convolutions [63] are employed to reduce the number of parameters and computational complexity, making the model more efficient.
+SK(Satoshi Kondo, Satoshi Kasai, Kousuke Hirasawa)
+Figure 9: Training de-artifact network and segmentation network. (Team SK)
+Team SK applies a two-step regime for the challenge.
+They use two networks: a de-artifact network and a segmentation network.
+The left subfigure of Figure 9 shows the de-artifact network.
+They add one of the three artifacts: brightness, smoke, and blood to the input image for training.
+The brightness artifacts are realized by reducing the brightness and contrast in a predefined range.
+The smoke artifacts are realized by alpha-blending an input image and an artificial smoke image.
+The blood artifacts are realized by adding red-colored ellipses to the input image.
+The de-artifact network is an encoder-decoder architecture like UNet [16].
+The encoder part is replaced by ConvNeXt-v2-base [64].
+The loss function is the L1 loss between the input image and the artifact-reduced image.
+The right subfigure of Figure 9 shows the segmentation network.
+The artifact addition block has the same functionality as the one when the de-artifact network is trained.
+For robustness, artifact addition is applied to randomly selected 75 % of the training datasets.
+The de-artifact network trained in the first step is frozen, and only the segmentation network is trained.
+The architecture of the segmentation network is the same as the de-artifact network.
+The loss function is a weighted summation of Dice loss, cross-entropy loss, and Hausdorff distance loss.
+Tailor(Sita Tailor, Ricardo Sanchez-Matilla, Imanol Luengo):
+Figure 10: Overview of the model and training. (Team Tailor)
+Team Tailor applies both surgical-specific and general augmentations to better simulate real surgical conditions.
+The surgical-specific augmentations included motion blur, low lighting, blood, and smoke, mimicking non-adversarial corruptions typical in operating environments.
+Additionally, general augmentations like affine transformations, flips, and random resized cropping are applied to introduce further variability.
+All augmentations were implemented using functions from the Albumentations library [65].
+They focus on transformer-based models for robust surgical tool segmentation, selecting the SwinV2 Transformer (SwinV2 small) [66] for its superior performance with segmentation tasks.
+They modify the model by replacing the classification head with a segmentation head, incorporating a 1x1 convolutional layer followed by upsampling to generate segmentation maps.
+An overview of the framework is displayed in Figure 10.
+FightTumor(Tianhao Fu, Jun Ma, Bo Wang)
+Team FightTumor’s approach uses pre-trained UNet [16] with AutoAugment [34].
+They customize augmentations in Albumentations [65].
+For smoke, the randomized smoke image is merged on top of the original image with an opacity of 1.
+They also generate a circle flare in which the border is blurred.
+ICVS-2AI(Marcos Fernández-Rodríguez, Estevão Lima, João L. Vilaça):
+For Team ICVS-2AI, a CycleGAN is trained with default parameters [67] and unpaired images, chosen from a private laparoscopic dataset, with 1496 clear and 220 containing smoke artifacts.
+To learn the “clear to smoke” style, they use the trained model with the SegSTRONG-C dataset, generate smoke versions of the original images, and use the original ground truth for the new images.
+Finally, the nnUNet network is trained using standard parameters described in their paper [68], with the increased dataset, using a combined dice and top-k loss.
+AIMI(Negin Ghamsarian, Klaus Schoeffmann, Raphael Sznitman)
+Team AIMI explore three architectures ReCalNet [69], AdaptNet [70], and DeepPyramid+ [71, 72].
+After thorough experiments, they finally apply DeepPyramid+ [71], which features two main modules: Pyramid View Fusion (PVF) and Deformable Pyramid Reception (DPR) as their architecture.
+The PVF module enhances pixel-level information representation by simulating the human visual system, offering a global-to-local perspective to capture relevant details around each pixel.
+The DPR module employs dilated deformable convolutions to facilitate shape- and scale-adaptive feature extraction, thereby improving segmentation accuracy for diverse and deformable classes within medical images.
+They apply common data augmentations like random cropping, color jittering, grayscale transformation, and Gaussian blur during training for better robustness.
+Data availability:
+The challenge data is released publicly and can be accessed at https://segstrongc.cs.jhu.edu/
+Code availability:
+The code is made available at https://github.com/hding2455/CaRTS/.
+Acknowledgement
 This Challenge is supported by the collaborative research agreement with the MultiScale Medical Robotics Center at The Chinese University of Hong Kong, Intuitive Surgical Inc, and Johns Hopkins University internal funds.
-Figure 3: Baseline results. The white area means true positive, the red area means false negative, and the orange area means false positive.